OpenStudy (anonymous):

If h(x) = 2x + 3, find the value of x if h(x) = 9. 21 3 6 9

OpenStudy (anonymous):

2x + 3 = 9 2x = 6 x = 3 plug it back into the function to check it h(3) = 2(3) + 3 = 6 + 3 = 9